# PatternKV: Flattening KV Representation Expands Quantization Headroom

### v1

@@ Abstract
KV cache in autoregressive LLMs eliminates redundant recomputation but has emerged as the dominant memory and bandwidth bottleneck during inference, notably with long contexts and test-time scaling. KV quantization is a key lever for reducing cache cost, but accuracy drops sharply as the native KV distribution lacks flatness and thus maintains a wide quantization range. Prior work focuses on isolating outliers, which caps their error but fails to flatten the overall distribution, leaving performance fragile under low-bit settings. In this work, we show that the K cache maintains a stable structure that evolves gradually with context, while the V cache carries latent semantic regularities. Building on these insights, we propose PatternKV, a pattern-aligned residual quantization scheme. It mines representative pattern vectors online, aligns each KV vector to its nearest pattern, and quantizes only the residual. This reshaping of the KV distribution flattens the quantization target and narrows its range, thereby improving the fidelity of low-bit KV quantization. Across long-context and test-time scaling settings on multiple backbones, PatternKV delivers consistent 2-bit gains, with a 0.08% average 4-bit drop relative to FP16, improves test-time scaling accuracy by 10% on average, and raises throughput by 1.4× while supporting 1.25× larger batches.
@@ 1 Introduction
Quantization (Ashkboos et al., 2024; Frantar et al., 2022) is a widely adopted approach for KV cache compression, reducing memory footprint via lower-bit KV representations. The effectiveness of KV quantization largely depends on the flatness of the vector distribution: flatter distributions yield a narrower quantization range and preserve higher precision under limited bit widths. In pursuit of this, Hooper et al. (2024); Kang et al. (2024); Su et al. (2025) handle outliers by storing them with original precision, separated from the main KV distribution to minimize their impact on quantization. Meanwhile, Liu et al. (2024b) confines outlier-induced quantization error by quantizing key cache per-channel, ensuring the error remains within individual channels. However, these methods are primarily limited to protecting outliers rather than flattening the entire distribution.
In contrast, we tackle the root cause of quantization inefficiency by reshaping the entire KV distribution. Guided by a variance–decomposition perspective, we mine common patterns in KV caches, align each vector to its nearest pattern, and quantize only the residuals. This distribution-wide treatment flattens the quantization target, yielding narrower ranges and substantially reducing error under low-bit settings.
Figure 1: The left figure illustrates the original distribution of the KV vectors, while the right figure depicts the distribution of the residuals obtained after aligning the original vectors with the corresponding pattern vectors. Each pattern vector is the centroid of its cluster.
Specifically, our analysis of KV caches reveals exploitable regularities: the K cache maintains a stable structure but will evolve gradually with context, while the V cache exhibits latent semantic patterns. These findings indicate that pattern information can be reliably mined online without calibration corpora or additional tuning. Building on this, we employ clustering to extract representative pattern vectors that capture such common structure. During inference, each KV vector is aligned to its nearest pattern vector and transformed into a residual for quantization, resulting in a markedly flatter distribution. To accommodate the gradual evolution of KV distributions over decoding, we further introduce new pattern vectors on the fly, adaptively tracking shifts and maintaining quantization fidelity.
@@
  • We evaluate our method against strong baselines across diverse tasks and backbone models. In the long-context setting, our approach achieves consistent gains at 2-bit while limiting the 4-bit average drop relative to FP16 to just 0.08%. Under test-time scaling, our method achieves a 10% average improvement. In addition, our method achieves a 1.4× throughput increase and supports a 1.25× larger batch size.
@@ 2.1 A Variance Decomposition View of KV Quantization
The law of total variance (Blitzstein & Hwang, 2019) is widely used for analyzing variance reductions (Depeweg et al., 2018; Lakshminarayanan et al., 2017). It states that, given a partition of the data into groups, the total variance can be decomposed into two components: an intra-group term and an inter-group term. To apply this principle in the KV setting, we can introduce a set of representative pattern vectors $M$ that partition the collection of KV vectors into different clusters. Under this view, the total variance of KV vectors $Z$ decomposes as
@@ 2.2.1 Origins and Evolution of K Cache Patterns
Prior work identifies outlier distributions in the K cache (Liu et al., 2024b; Hooper et al., 2024), and we extend this line of evidence with a systematic robustness analysis (Appendix B), which shows that a fixed model’s K cache maintains a stable structure attributable to internal linear mappings and nonlinear activations rather than any particular prompt. To probe this origin, we run an input–decoupling experiment: for each token, we compare the K cache distribution when propagating only the token embedding to that obtained from the full hidden state carrying context. As shown in Fig. 2, outlier channels already appear with embedding-only input, adding context chiefly inflates overall magnitude and dynamic range while leaving the structural pattern intact. The invariance of this pattern across inputs indicates that reliable pattern estimates can be obtained directly from the observed activations, without heavy dependence on corpus-specific calibration. We conclude:
Figure 2: Channel-wise mean absolute value distributions. Left: embedding-only injection; Right: full-input injection. Outlier channels are already evident under embedding-only input, and the full input further enlarges the range and extremes. Additional figures are provided in Appendix C.
@@
Figure 3: (a) t-SNE visualization of the of K-cache distributions across attention heads along a single inference trajectory. (b) Illustration of the degree of alignment between V cache clusters and semantic categories. Additional figures are provided in Appendix C.
Building on Insight 1, we analyze how the evolving context reshapes the K cache distribution during decoding. We sample K vectors along a single inference trajectory and visualize them per attention head using t-SNE. As shown in Fig. 3(a), the K distribution drifts smoothly across decoding steps rather than exhibiting abrupt jumps, and each head follows a distinct trajectory. This behavior is consistent with rotary positional embeddings, which inject relative position after Q and K are formed. Notably, although the marginal distribution evolves, short-range geometry remains stable: nearby tokens along the sequence tend to inhabit similar regions. This local consistency makes it natural to ground pattern estimates in the immediate neighborhood along the trajectory, where local similarity is highest. Hence,
@@ 2.2.2 Analysis of Latent Pattern in V Cache
We then aggregate $C_{t}$ across layers and heads to assess within-cluster cohesion. As shown in Fig. 3(b), shallow and deep layers exhibit strong alignment between tokens and cluster assignments, supporting our hypothesis. In the middle layers, the same token spreads across multiple clusters, indicating weaker coupling between V representations and semantics, which hampers the extraction of common structure. Therefore,
Insight 3 The V cache generally exhibits latent semantic patterns, with the association remaining strong in most layers and attenuating in some middle layers.
@@ 3 Method
In light of the previous analysis, we propose PatternKV, a residual quantization pipeline based on pattern alignment, as illustrated in Fig. 4. In the prefill stage, we select pattern vectors online via clustering to minimize within-pattern variance (Insight 1). In the decode stage, we update the pattern vector to the Chebyshev center to adaptively track the distribution’s gradual evolution (Insight 2). For pattern utilization, we assign each KV vector to a pattern using the min–max distance and quantize only the residual, which flattens the target distribution and contracts its dynamic range. For the V cache, where semantic alignment is weaker in intermediate layers (Insight 3), we further incorporate an adaptive threshold so that flattening provably incurs error no greater than raw quantization. Besides, we provide one theoretical guarantee for the method in Appendix D.
@@ 3.1 Pattern Mining
$\min_{\mathcal{P}_{h}=P_{1},\ldots,P_{k}}\;\sum_{j=1}^{k}\;\sum_{\bm{x_{i}}\in P_{j}}\bigl\|\bm{x_{i}}-\bm{\overline{x}_{P_{j}}}\bigr\|_{2}^{2}$ (4)
Let $P_{k}$ denote the $k$-th pattern cluster. We optimize the objective using KMeans (McQueen, 1967) under the Euclidean metric and take the centroid of each cluster as its pattern vector. For the $h$-th attention head, the resulting set of pattern vectors is $\mathcal{M}_{h}=\{M_{1},\ldots,M_{k}\}$.
Since our objective coincides with the K-means objective, the partition returned at convergence is a local minimizer of the within-pattern variance.
@@
Guided by Insight 2, we update per-head pattern vectors during decoding to adaptively track the distribution’s gradual evolution. Instead of arithmetic means, we use Chebyshev centers computed over each group of KV vectors, which minimize the local quantization range and provide stronger robustness to outliers, thereby aligning better with the asymmetric quantization objective.
Specifically, we use the quantization group window $G_{\text{pattern}}$ to generate new pattern vectors. For the pattern vector $M_{h}^{new}$ of the $h$-th attention head within this window, we have:
@@ 3.2 Pattern Utilization
Specifically, we adopt the min–max distance for pattern selection, defined for a vector $\bm{x}$ and a candidate pattern $\bm{m}$ as $d_{mm}(\bm{x},\bm{m})=\max_{i}(x_{i}-m_{i})-\min_{j}(x_{j}-m_{j})$.
@@ 3.3 Flattening-Sensitive Adaptive Threshold for V Pattern Utilization
$H_{0}:\ \mathbb{E}[D]\leq 0$
$\mathbb{E}[D]=\frac{\Delta_{raw}^{2}-\Delta_{flat}^{2}}{12},\qquad\mathrm{Var}(D)=\frac{\Delta_{raw}^{4}+\Delta_{flat}^{4}}{180\,d}$ (8)
$\frac{\mathbb{E}[D]-0}{\sqrt{\mathrm{Var}(D)}}\geq z_{1-\alpha}$ (9)
$1-\rho^{2}\;\geq\;\frac{2\,z_{1-\alpha}}{\sqrt{5\,d}}\;\sqrt{1+\rho^{4}}\quad\Longleftrightarrow\quad\rho\;\leq\;\rho_{*}(d,\alpha)$ (10)
4 Experiments
4.1 Settings
Benchmarks
As long contexts and test-time scaling commonly render the KV cache the dominant memory and bandwidth bottleneck during inference, we structure our evaluation into two categories. For the long-input setting, we use the full LongBench (Bai et al., 2024) benchmark, which offers multiple evaluation dimensions with task-specific metrics. LongBench details appear in Appendix E. For reasoning, we consider GSM8K (Cobbe et al., 2021), AIME (Balunović et al., 2025), and AMC (Li et al., 2024a). GSM8K probes the impact of quantization on chain-of-thought capability, and AIME and AMC evaluate performance under long chain-of-thought scenarios.
Models
Baselines
Because our method is an online algorithm that requires no offline calibration set, we compare it against online quantization baselines: KIVI (Liu et al., 2024b), ZipCache (He et al., 2024), SKVQ (Duanmu et al., 2024) and OTT (Su et al., 2025). Detailed experimental settings for the baseline methods are provided in Appendix F.
Quantization Settings
4.2 Main Results
Results on LongBench
Table 1: Overall LongBench results at 2-bit precision. The best and second-best in every column are marked in bold and underline, respectively. See Appendix G for the 4-bit precision results.
@@ 4.2 Main Results
We evaluate on all 21 datasets of LongBench, focusing on two quantization precisions: INT2 and INT4. The 2-bit results in Table 1 demonstrate that our approach achieves robust gains over competitive baselines despite the extreme precision constraint. While some baselines achieve notable improvements on code-related tasks yet fail to generalize to other categories, our method provides stable and consistent improvements across task types. Results for INT4 setting are provided in Appendix G.
Results on Long-CoT Settings
Test-time scaling improves LLM reasoning through depth-oriented expansion and breadth-oriented expansion, which yields long outputs and substantially increases KV cache usage. To this end, we evaluate models that can generate long chain-of-thought rationales on challenging mathematical benchmarks. For each problem, we generate eight independent responses and report $Avg@8$ (per-sample accuracy averaged over the eight responses) and $Maj@8$ (problem-level accuracy under majority voting across the eight responses). Table 2 reports the INT2 results: prior methods degrade markedly, whereas our method achieves an average 10% improvement. We also evaluate under the INT4 setting, detailed results are provided in Appendix H.
Table 2: Overall Results on the Long-CoT Benchmark at 2-bit precision. See Appendix H for the 4-bit precision results.
@@
Figure 5: GSM8K accuracy under zero-shot CoT on Llama-3.1-8B-Instruct.

### v2

@@ Abstract
KV cache in autoregressive LLMs eliminates redundant recomputation but has emerged as the dominant memory and bandwidth bottleneck during inference, notably with long contexts and test-time scaling. KV quantization is a key lever for reducing cache cost, but accuracy drops sharply as the native KV distribution lacks flatness and thus maintains a wide quantization range. Prior work focuses on isolating outliers, which caps their error but fails to flatten the overall distribution, leaving performance fragile under low-bit settings. In this work, we show that the K cache maintains a stable, context-evolving structure, while the V cache carries latent semantic regularities, with both contributing to the organization of vectors into shared patterns. Building on these insights, we propose PatternKV, a pattern-aligned residual quantization scheme. It mines representative pattern vectors online, aligns each KV vector to its nearest pattern, and quantizes only the residual. This reshaping of the KV distribution flattens the quantization target and narrows its range, thereby improving the fidelity of low-bit KV quantization. Across long-context and test-time scaling settings on multiple backbones, PatternKV delivers consistent 2-bit gains, with a 0.08% average 4-bit drop relative to FP16, improves test-time scaling accuracy by 10% on average, and raises throughput by 1.5× while supporting 1.25× larger batches.
Machine Learning, ICML
@@ 1 Introduction
Quantization (Ashkboos et al., 2024; Frantar et al., 2022) is a widely adopted approach for KV cache compression, reducing memory footprint via lower-bit KV representations. The effectiveness of KV quantization largely depends on the flatness of the vector distribution (see Figure 1(a) and Section 2.1): flatter distributions yield a narrower quantization range and preserve higher precision under limited bit widths. In pursuit of this, Hooper et al. (2024); Kang et al. (2024); Su et al. (2025) handle outliers by storing them with original precision, separated from the main KV distribution to minimize their impact on quantization. Meanwhile, Liu et al. (2024b) confines outlier-induced quantization error by quantizing key cache per-channel, ensuring the error remains within individual channels. However, these methods are primarily limited to protecting outliers rather than flattening the entire distribution.
In contrast, we tackle the root cause of quantization inefficiency by reshaping the entire KV distribution (Figure 1(b)). Guided by a variance–decomposition perspective, we mine common patterns in KV caches, align each vector to its nearest pattern, and quantize only the residuals. This distribution-wide treatment flattens the quantization target, yielding a narrower range and substantially reducing error under low-bit settings.
Figure 1: (a) Impact of vector-distribution flatness on KV quantization. Flatter value distributions (top) concentrate samples within a narrower quantization range after zero-point shifting (Xzp)(X-zp), thereby improving the utilization of discrete levels and preserving higher precision at constrained bit widths. In contrast, peaked distributions with outliers (bottom) broaden the required range, leaving many levels underutilized and degrading precision. (b) The left figure illustrates the original distribution of the KV vectors, while the right figure depicts the distribution of the residuals obtained after aligning the original vectors with the corresponding pattern vectors. Each pattern vector is the centroid of its cluster.
Specifically, our analysis of KV caches reveals exploitable regularities: the K cache maintains a stable structure that will evolve gradually with context, and the V cache exhibits latent semantic patterns that induce token-consistent structure in the V vectors. These findings indicate that pattern information can be reliably mined online without calibration corpora or additional tuning. Building on this, we employ clustering to extract representative pattern vectors that capture such common structure. During inference, each KV vector is aligned to its nearest pattern vector and transformed into a residual for quantization, resulting in a markedly flatter distribution. To accommodate the gradual evolution of KV distributions over decoding, we further introduce new pattern vectors on the fly, adaptively tracking shifts and maintaining quantization fidelity.
@@
  • We evaluate our method against strong baselines across diverse tasks and backbone models. In the long-context setting, our approach achieves consistent gains at 2-bit while maintaining near-lossless 4-bit accuracy (only 0.08% lower than FP16 on average). Under test-time scaling, our method achieves a 10% average improvement. In addition, our method achieves a 1.5× throughput increase and supports a 1.25× larger batch size.
@@ 2 Motivations
Figure 2: Channel-wise mean absolute value distributions. Left: embedding-only injection; Right: full-input injection. Outlier channels are already evident under embedding-only input, and the full input further enlarges the range and extremes. Additional figures are provided in Appendix B.
@@ 2.1 A Variance Decomposition View of KV Quantization
The law of total variance (Blitzstein and Hwang, 2019) is widely used for analyzing variance reductions (Depeweg et al., 2018; Lakshminarayanan et al., 2017). It states that, given a partition of the data into groups, the total variance can be decomposed into two components: an intra-group term and an inter-group term. To apply this principle in the KV setting, we can introduce a set of representative pattern vectors $M$ that partition the collection of KV vectors into different clusters. Under this view, the total variance of KV vectors $Z$ decomposes as
@@ 2.2.1 Origins and Evolution of K Cache Patterns
Prior work identifies outlier distributions in the K cache (Liu et al., 2024b; Hooper et al., 2024), and we extend this line of evidence with a systematic robustness analysis (Appendix A), which shows that a fixed model’s K cache maintains a stable structure attributable to internal linear mappings and nonlinear activations rather than any particular prompt. To probe this origin, we run an input–decoupling experiment: for each token, we compare the K cache distribution when propagating only the token embedding to that obtained from the full hidden state carrying context. As shown in Figure 2, outlier channels already appear with embedding-only input, adding context chiefly inflates overall magnitude and dynamic range while leaving the structural pattern intact. The invariance of this pattern across inputs indicates that reliable pattern estimates can be obtained directly from the observed activations, without heavy dependence on corpus-specific calibration. We conclude:
@@
Figure 3: (a) t-SNE visualization of the of K-cache distributions across attention heads along a single inference trajectory. (b) Illustration of the degree of alignment between V cache clusters and semantic categories. Additional figures are provided in Appendix B.
Building on Insight 1, we analyze how the evolving context reshapes the K cache distribution during decoding. We sample K vectors along a single inference trajectory and visualize them per attention head using t-SNE. As shown in Figure 3(a), the K distribution drifts smoothly across decoding steps rather than exhibiting abrupt jumps, and each head follows a distinct trajectory. This behavior is consistent with rotary positional embeddings, which inject relative position after Q and K are formed. Notably, although the marginal distribution evolves, short-range geometry remains stable: nearby tokens along the sequence tend to inhabit similar regions. This local consistency makes it natural to ground pattern estimates in the immediate neighborhood along the trajectory, where local similarity is highest. Hence,
@@ 2.2.2 Analysis of Latent Pattern in V Cache
We then aggregate $C_{t}$ across layers and heads to assess within-cluster cohesion. As shown in Figure 3(b), shallow and deep layers exhibit strong alignment between tokens and cluster assignments, supporting our hypothesis. In the middle layers, the same token spreads across multiple clusters, indicating weaker coupling between V representations and semantics, which hampers the extraction of common structure. Therefore,
Insight 3 The V cache generally exhibits latent semantic patterns, with token–cluster association remaining strong in most layers and attenuating in some middle layers.
@@ 3 Method
In light of the previous analysis, we propose PatternKV, a residual quantization pipeline based on pattern alignment, as illustrated in Figure 4. In the prefill stage, we select pattern vectors online via clustering to minimize within-pattern variance (Insight 1). In the decode stage, we update the pattern vector to the Chebyshev center to adaptively track the distribution’s gradual evolution (Insight 2) and the potential emergence of new semantic patterns during generation (Insight 3). For pattern utilization, we assign each KV vector to a pattern using the min–max distance and quantize only the residual, which flattens the target distribution and contracts its dynamic range. For the V cache, where semantic alignment is weaker in intermediate layers (Insight 3), we further incorporate an adaptive threshold so that flattening provably incurs error no greater than raw quantization. Besides, we provide one theoretical guarantee for the method in Appendix C.
@@ 3.1 Pattern Mining
$\min_{\mathcal{P}_{h}=P_{1},\ldots,P_{k}}\;\sum_{j=1}^{k}\;\sum_{\boldsymbol{x_{i}}\in P_{j}}\bigl\|\boldsymbol{x_{i}}-\boldsymbol{\overline{x}_{P_{j}}}\bigr\|_{2}^{2}$ (4)
Let $P_{k}$ denote the $k$-th pattern cluster. We optimize the objective using KMeans (McQueen, 1967) under the Euclidean metric and take the centroid of each cluster as its pattern vector. For the $h$-th attention head, the resulting set of pattern vectors is $\mathcal{M}_{h}=\{M_{1},\ldots,M_{k}\}$. Since our objective coincides with the K-means objective, the partition returned at convergence is a local minimizer of the within-pattern variance.
@@
Guided by Insight 2, we update per-head pattern vectors during decoding to adaptively track the distribution’s gradual evolution. Instead of arithmetic means, we use Chebyshev centers computed over each group of KV vectors, which minimize the local quantization range and provide stronger robustness to outliers, thereby aligning better with the asymmetric quantization objective. Specifically, we use the quantization group window $G_{\text{pattern}}$ to generate new pattern vectors. For the pattern vector $M_{h}^{new}$ of the $h$-th attention head within this window, we have:
@@ 3.2 Pattern Utilization
Specifically, we adopt the min–max distance for pattern selection, defined for a vector $\boldsymbol{x}$ and a candidate pattern $\boldsymbol{m}$ as $d_{mm}(\boldsymbol{x},\boldsymbol{m})=\max_{i}(x_{i}-m_{i})-\min_{j}(x_{j}-m_{j})$.
@@ 3.3 Flattening-Sensitive Adaptive Threshold for V Pattern Utilization
Table 1: Overall LongBench results at 2-bit precision. The best and second-best in every column are marked in bold and underline, respectively. See Appendix F for the 4-bit precision results.
@@
$H_{0}:\ \mathbb{E}[D]\leq 0$
$\mathbb{E}[D]=\frac{\Delta_{raw}^{2}-\Delta_{flat}^{2}}{12},\mathrm{Var}(D)=\frac{\Delta_{raw}^{4}+\Delta_{flat}^{4}}{180\,d}$ (8)
$\frac{\mathbb{E}[D]-0}{\sqrt{\mathrm{Var}(D)}}\geq z_{1-\alpha}$ (9)
$1-\rho^{2}\;\geq\;\frac{2\,z_{1-\alpha}}{\sqrt{5\,d}}\;\sqrt{1+\rho^{4}}\;\Longleftrightarrow\;\rho\;\leq\;\rho_{*}(d,\alpha)$ (10)
3.4 System Optimization
Prefill-stage pattern extraction
Because each attention head requires its own set of pattern vectors, we implement a fully GPU-parallel KMeans algorithm to minimize the clustering latency during inference.
Decode-stage KV reconstruction
Reconstructing KV vectors during decoding introduces substantial overhead, so we design two customized CUDA kernels to reduce the cost. The first kernel fuses pattern index restoration for the K cache, dequantization, and the QK matrix multiplication into a single operator. The second kernel is designed to fuse dequantization, pattern index restoration, and attention weight application over V into one operator. Together, these fused kernels significantly reduce decode-stage latency compared with a pure PyTorch implementation.
Following the above system optimizations, we achieve an approximately 15% increase in peak throughput on a single GPU and a 10% reduction in per-sample inference latency, further demonstrating the effectiveness of our optimizations. [^1] Our code is available in our GitHub repository: https://github.com/HCOOOH/PatternKV.
4 Experiments
Table 2: Overall Results on the Long-CoT Benchmark at 2-bit precision. See Appendix G for the 4-bit precision results.
@@ 4 Experiments
4.1 Settings
Benchmarks
As long contexts and test-time scaling commonly render the KV cache the dominant memory and bandwidth bottleneck during inference, we structure our evaluation into two categories. For the long-input setting, we use the full LongBench (Bai et al., 2024) benchmark, which offers multiple evaluation dimensions with task-specific metrics. LongBench details appear in Appendix D. For reasoning, we consider GSM8K (Cobbe et al., 2021), AIME (Balunović et al., 2025), and AMC (Li et al., 2024a). GSM8K probes the impact of quantization on chain-of-thought capability, and AIME and AMC evaluate performance under long chain-of-thought scenarios.
Models
Baselines
Because our method is an online algorithm that requires no offline calibration set, we compare it against online quantization baselines: KIVI (Liu et al., 2024b), ZipCache (He et al., 2024), SKVQ (Duanmu et al., 2024) and OTT (Su et al., 2025). Detailed experimental settings for the baseline methods are provided in Appendix E.
Quantization Settings
4.2 Main Results
Results on LongBench
We evaluate on all 21 datasets of LongBench, focusing on two quantization precisions: INT2 and INT4. The 2-bit results in Table 1 demonstrate that our approach achieves robust gains over competitive baselines despite the extreme precision constraint. While some baselines achieve notable improvements on code-related tasks yet fail to generalize to other categories, our method provides stable and consistent improvements across task types. Results for INT4 setting are provided in Appendix F.
Results on Long-CoT Settings
Test-time scaling improves LLM reasoning through depth-oriented expansion and breadth-oriented expansion, which yields long outputs and substantially increases KV cache usage. To this end, we evaluate models that can generate long chain-of-thought rationales on challenging mathematical benchmarks. For each problem, we generate eight independent responses and report $Avg@8$ (per-sample accuracy averaged over the eight responses) and $Maj@8$ (problem-level accuracy under majority voting across the eight responses). Table 2 reports the INT2 results: prior methods degrade markedly, whereas our method achieves an average 10% improvement. We also evaluate under the INT4 setting, detailed results are provided in Appendix G.